Calculate the energy values of the electron in the second orbit of the hydrogen atom. What will be the radius of this orbit
1
Expert's answer
2019-03-18T02:30:03-0400

The expression for energy values and the radius of a certain atomic orbit, according to the Bohr model, is: Energy(E) = -13.6 x Z2/n2 eV/atom Radius(r) = 0.529 x Z2/n Ao 1eV=1.6x 10−19J ; For second orbit of hydrogen atom, n=2 and Z=1, so, E = -13.6 x 1/4 = -3.4 eV r = 0.529 x 1/2 = 0.2645 Ao